OpenStudy (anonymous):

Pretty lost on this problem! Help needed! Q: Suppose that 3 less than or equal to f^(x) is less than or equal to 5 for all values of x? Show that 18 is less than or equal to f(8)- f(2) is less than or equal to 30.

OpenStudy (anonymous):

I dont really know where to start..

OpenStudy (dan815):

Show that 18 is less than or equal to f(8)- f(2) is less than or equal to 30. what is thsi part saying

OpenStudy (anonymous):

Here lemme right it out: Suppose that \[3\le f'x \le 5 \]

OpenStudy (dan815):

oh xD thats a derivative haha

OpenStudy (anonymous):

and show that \[18 \le f(8) - f(12) \le 30 \]

OpenStudy (dan815):

okk

OpenStudy (anonymous):

lol yea

OpenStudy (dan815):

basically, lets consider the max slope and min slope

OpenStudy (dan815):

f'(x) has a min of 3 and max of 5 slope, what can we say about the distance change from f(8) to f(12) then

OpenStudy (dan815):

i see you wrote f(2) up in your question then f(12) rewritten which one is it exactly?

OpenStudy (anonymous):

its f(2), sorry about that

OpenStudy (dan815):

okay so

OpenStudy (dan815):

do you have any ideas so far?

OpenStudy (anonymous):

how would you solve for what f(8) and f(2) is if they dont even give you what the orignal function is?

OpenStudy (dan815):

we dont know f(8) and f(2) but we can see the max and min difference

OpenStudy (dan815):